# Robust Locomotion via Zero-order Stochastic Nonlinear Model Predictive Control with Guard Saltation Matrix

### v1

@@ Abstract
This paper presents a stochastic/robust nonlinear model predictive control (NMPC) to enhance the robustness of legged locomotion against contact uncertainties. We integrate the contact uncertainties into the covariance propagation of stochastic/robust NMPC framework by leveraging the guard saltation matrix and an extended Kalman filter-like covariance update. We achieve fast stochastic/robust NMPC computation by utilizing the zero-order stochastic/robust NMPC algorithm with additional improvements in computational efficiency concerning the feedback gains. We conducted numerical experiments and demonstrate that the proposed method can accurately forecast future state covariance and generate trajectories that satisfies constraints even in the presence of the contact uncertainties. Hardware experiments on the perceptive locomotion of a wheeled-legged robot were also carried out, validating the feasibility of the proposed method in a real-world system with limited on-board computation.
@@ II HYBRID SYSTEM AND GUARD SALTATION MATRIX
The original saltation matrix [11] describes the first-order sensitivity of $x(t_{(m,m+1)}^{+})$ with respect to $x(t_{(m,m+1)}^{-})$ in cases where (2) is one-dimensional, i.e., (2) is a scalar-valued function. Specifically, the saltation matrix for the switch from mode $m$ to $m+1$ is defined as:
@@ III-A Problem Settings
where $\Gamma_{i}\in\mathbb{R}^{n_{x}\times n_{w}}$ is a constant matrix, and $m$ is the appropriate mode index. We also consider disturbances in the state jump:
@@
where $A_{i}:=\nabla_{x}F_{i}(t_{i},\bar{x}_{i},u_{i},)$, $B_{i}:=\nabla_{u}F_{i}(t_{i},\bar{x}_{i},u_{i},0)$, $\Gamma_{i}:=\nabla_{w}F_{i}(t_{i},\bar{x}_{i},u_{i},0)$, $W_{i}\in\mathbb{R}^{n_{w}\times n_{w}}$ is a symmetric positive semi-definite matrix, and $K_{i}\in\mathbb{R}^{n_{u}\times n_{x}}$ is a state feedback gain. Equation (III-A) refers to closed-loop covariance propagation with an appropriate feedback gain $K_{i}$, and reduces to open-loop propagation with $K_{i}=O$. Closed-loop propagation can mitigate the overgrowth of state covariance and result in less conservative constraints [5], but requires optimization of the feedback gain $K_{i}$. In robust MPC, $W_{i}$ describes an ellipsoid containing the process noise $w_{i}$ [4, 5]. In stochastic MPC, $W_{i}$ describes the noise distribution covariance, such as $w_{i}\sim\mathcal{N}(0,W_{i})$ [6].
@@ III-B2 A posteriori covariance update using output-feedback MPC
The update (III-B1) is an open-loop covariance update and may lead to overgrowth of the covariance. To address this, we apply an a posteriori covariance update from output-feedback stochastic MPC [3]. For the stochastic process whose blief dynamics are (7) and (8), the guard condition (2), i.e., the contact positions, can be regarded an output equation, as is well-known in the EKF-based state estimation for the legged robots [14]. Treating the guard condition (2) as an output equation, we apply the a posteriori covariance update law proposed in [3], resulting in:
@@ III-C Stochastic/Robust Optimal Control Problem
$\min_{\begin{subarray}{c}\bar{x}_{0},...,\bar{x}_{N},\\ \bar{u}_{0},...,\bar{u}_{N-1},\\ P_{0},...,P_{N}\end{subarray}}$ $J:=V_{N}(\bar{x}_{N})+\sum_{i\in I}l_{i}(\bar{x}_{i},\bar{u}_{i})+\sum_{j\in J}l_{j}(\bar{x}_{i})$ (14d)
${\rm s.t.}\;\;$ $\bar{x}_{0}=\bar{x},$ (14e)
$\bar{x}_{i+1}=F_{i}(t_{i},\bar{x}_{i},\bar{u}_{i},0),\;\;i\in I,$ (14f)
$\bar{x}_{j+1}=F_{j}(t_{j},\bar{x}_{j},0),\;\;j\in J,$ (14g)
$P_{i+1}=\Phi_{i}(P_{i},K_{i},\bar{x}_{i},\bar{u}_{i}),\;\;i\in I,$ (14h)
$P_{j+1}=\Phi_{j}(P_{j},\bar{x}_{j}),\;\;j\in J,$ (14i)
$g_{j}(\bar{x}_{j})=0,\;\;j\in J,$ (14j)
$h_{i}(\bar{x}_{i},\bar{u}_{i})+\beta_{i}(P_{i},K_{i},\bar{x}_{i},\bar{u}_{i})\geq 0,\;i\in I,$ (14k)
$h_{j}(\bar{x}_{j})+\beta_{j}(P_{j},\bar{x}_{j})\geq 0,\;j\in J,j=N,$ (14l)
@@ IV-A Zero-Order Algorithm for Covariance Propagation
The zero-order algorithm of stochastic/robust NMPC [4, 5, 3] simplifies the Newton-type iteration for the OCP (III-C) by disregarding the sensitivities of the covariance propagation equations (14h) and (14i) as well as the backoff terms $\beta(\cdot)$. The computational burden per Newton-type iteration is competitive with the nominal NMPC.
A limitation of the zero-order algorithm is that its convergence rate may slow down when the process noise covariance $W_{i}$ is substantial, as the backoff terms have a large effect on the convergence. We empirically found that for efficient optimization, $W_{i}$ should be set to moderate the covariance propagation rather than accurately reflecting the process noise covariance of the actual system.
@@ IV-B Interior Point Method and Feedback Gain
We employ the nonlinear primal-dual interior point method to handle hard inequality constraints [15]. It should be noted that state-only inequality constraints need to be treated as soft constraints such as [16] to prevent the backoff term (III-C) from becoming excessively large, which leads to no feasible solution satisfying the tightened constraints (14k)–(14l) [4]. In practice, we also might have to clip the maximum values of the backoff terms to avoid ill-conditioned optimization problems.
@@ IV-C Algorithm Summary
6:     Propagate the state covariances by (14h) and (14i).
@@ V-A Full-Centroidal Formulations
Note that, although the reset map is the identity, the saltation matrix (II) is not the identity, and the guard saltation matrix (5) is not zero, due to the different values of $f_{m}^{-}$ and $f_{m+1}^{+}$. The guard condition consists of the distances between the swing feet and their targeted contact surfaces. Given a configuration $q$, the guard condition and its derivatives can be computed using forward kinematics and its Jacobian
@@ V-B Comparison in Covariance Prediction
(a) A priori and a posteriori covariance updates (III-B1)–(13)
(b) A priori covariance update (III-B1)
@@
Fig. 3 illustrates the predicted base pose covariances using the three different methods. The state covariance is a $(12+n_{J})\times(12+n_{J})$ matrix, and the plots show the diagonal elements corresponding to the base pose dimensions. The comparison between methods (a) and (b) reveals that relying solely on the a priori covariance update (III-B1) leads to an overgrowth of the state covariance. This can lead to excessively large back-off terms (III-C) and (20), which in turn leads to a conservative solution or makes the numerical optimization challenging to solve. Conversely, the results from method (b) demonstrate that the combined use of a priori and a posteriori covariance updates effectively avoids the overgrowth of state covariance.
@@ V-C Simulation Comparison
Fig. 4 shows the success rates of the four methods in the two benchmarks. The proposed GS-SMPC outperformed the other methods in both scenarios, demonstrating its robustness against contact uncertainties. Figs. 57 display typical behaviors of Tachyon 3 and EVAL-03 during the experiments. As illustrated in Figs. 5 and 6, GS-SMPC successfully generated trajectories that avoided collisions while preventing joint torque satirations despite the discrepancies between expected and actual terrain heights. While SMPC also managed to avoid collisions, it encountered ill-conditioned optimization problems (e.g., too small step size in Fig. 6) caused by overly large backoff terms. Is should be noted that even we attempted SMPC with wide range of $\sigma$ in (23), but could not find successfull one. HMPC and MPC were unable to accurately predict constraint margins, leading to intense collisions following the terrain height mispredictions. Fig. 7 also demonstrated effectiveness of GS-SMPC in preventing self-collisions while enhancing the unilateral force constraints under terrain height disturbances. The robot frequently experienced self-collisions between the left and right legs due to the contact disturbances when using the other MPC methods. However, as suggested by the success rate for EVAL-03 in Fig. 4, the proposed method still encountered challenges in preventing the robot from falling down, e.g., when a swing foot collided with bumps or the terrain was too tilted.
(a) GS-SMPC
@@ V-D Hardware Experiments
Fig. 8 displays snapshots and plots of Tachyon 3’s perceptive locomotion using the proposed stochastic NMPC. Tachyon 3 adeptly navigated over a step and managed to ascend and descend a series of two steps via the proposed stochastic NMPC, which was implemented on the robot’s on-board PC. The average computational time of the stochastic NMPC was 29 ms while that of the nominal NMPC (i.e., NMPC without covariance propagation) was 26 ms measured on the on-board PC (CPU: Intel(R) Core(TM) i7-8850 H CPU @ 2.606 GHz). Thanks to the efficient algorithm of the stochastic NMPC outlined in Section IV, our stochastic/robust NMPC showed only a slight increase (approximately 3 ms in this instance) in computational time compared to the nominal NMPC.
@@ VI CONCLUSIONS
This paper introduced a stochastic/robust NMPC for robust legged locomotion against contact uncertainties. We leveraged the saltation matrices and output-feedback MPC technique in forecasting the future state covariance, of which the performance had been shown in the numerical studies. Furthermore, the proposed method could be implemented on a on-board PC by utilizing the efficient zero-order stochastic/robust NMPC algorithm that fully leverages the Riccati recursion to obtain the feedback gains.

### v2

# Robustifying Model-Based Locomotion by Zero-order Stochastic Nonlinear Model Predictive Control with Guard Saltation Matrix

@@ Abstract
This paper presents a stochastic/robust nonlinear model predictive control (NMPC) to enhance the robustness of model-based legged locomotion against contact uncertainties. We integrate the contact uncertainties into the covariance propagation of stochastic/robust NMPC framework by leveraging the guard saltation matrix and an extended Kalman filter-like covariance update. We achieve fast stochastic/robust NMPC computation by utilizing the zero-order algorithm with additional improvements in computational efficiency concerning the feedback gains. We conducted numerical experiments and demonstrate that the proposed method can accurately forecast future state covariance and generate trajectories that satisfies constraints even in the presence of the contact uncertainties. Hardware experiments on the perceptive locomotion of a wheeled-legged robot were also carried out, validating the feasibility of the proposed method in a real-world system with limited on-board computation.
@@ II HYBRID SYSTEM AND GUARD SALTATION MATRIX
The original saltation matrix [11] describes the first-order sensitivity of $x(t_{(m,m+1)}^{+})$ with respect to $x(t_{(m,m+1)}^{-})$ in cases where (2) is one-dimensional, i.e., (2) is a scalar-valued function. Specifically, the saltation matrix for the switch from mode $m$ to $m+1$ is defined as [11]:
@@ III-A Problem Settings
where $\Gamma_{i}\in\mathbb{R}^{n_{x}\times n_{w}}$ is a constant matrix, $m$ is the appropriate mode index, and $\Delta\tau>0$ is the time step. We also consider disturbances in the state jump:
@@
where $A_{i}:=\nabla_{x}F_{i}(t_{i},\bar{x}_{i},u_{i},0)$, $B_{i}:=\nabla_{u}F_{i}(t_{i},\bar{x}_{i},u_{i},0)$, $\Gamma_{i}:=\nabla_{w}F_{i}(t_{i},\bar{x}_{i},u_{i},0)$, $W_{i}\in\mathbb{R}^{n_{w}\times n_{w}}$ is a symmetric positive semi-definite matrix, and $K_{i}\in\mathbb{R}^{n_{u}\times n_{x}}$ is a state feedback gain. Equation (III-A) refers to closed-loop covariance propagation with an appropriate feedback gain $K_{i}$, and reduces to open-loop propagation with $K_{i}=O$. Closed-loop propagation can mitigate the overgrowth of state covariance and result in less conservative constraints [5], but requires optimization of the feedback gain $K_{i}$. In robust MPC, $W_{i}$ describes an ellipsoid containing the process noise $w_{i}$ [4, 5]. In stochastic MPC, $W_{i}$ describes the noise distribution covariance, such as $w_{i}\sim\mathcal{N}(0,W_{i})$ [6].
@@ III-B2 A posteriori covariance update using output-feedback MPC
The update (III-B1) is an open-loop covariance update and may lead to overgrowth of the covariance. To address this, we apply an a posteriori covariance update from output-feedback stochastic MPC [3]. For the stochastic process whose belief dynamics are (7) and (8), the guard condition (2), i.e., the contact positions, can be regarded an output equation, as is well-known in the EKF-based state estimation for the legged robots [14]. Treating the guard condition (2) as an output equation, we apply the a posteriori covariance update law proposed in [3], resulting in:
@@ III-C Stochastic/Robust Optimal Control Problem
$\min_{\begin{subarray}{c}\bar{x}_{0},...,\bar{x}_{N},\\ \bar{u}_{0},...,\bar{u}_{N-1},\\ P_{0},...,P_{N}\end{subarray}}$ $J:=V_{N}(\bar{x}_{N})+\sum_{i\in I}l_{i}(\bar{x}_{i},\bar{u}_{i})+\sum_{j\in J}l_{j}(\bar{x}_{i})$ (14a)
${\rm s.t.}\;\;$ $\bar{x}_{0}=\bar{x},$ (14b)
$\bar{x}_{i+1}=F_{i}(t_{i},\bar{x}_{i},\bar{u}_{i},0),\;\;i\in I,$ (14c)
$\bar{x}_{j+1}=F_{j}(t_{j},\bar{x}_{j},0),\;\;j\in J,$ (14d)
$P_{i+1}=\Phi_{i}(P_{i},K_{i},\bar{x}_{i},\bar{u}_{i}),\;\;i\in I,$ (14e)
$P_{j+1}=\Phi_{j}(P_{j},\bar{x}_{j}),\;\;j\in J,$ (14f)
$g_{j}(\bar{x}_{j})=0,\;\;j\in J,$ (14g)
$h_{i}(\bar{x}_{i},\bar{u}_{i})+\beta_{i}(P_{i},K_{i},\bar{x}_{i},\bar{u}_{i})\geq 0,\;i\in I,$ (14h)
$h_{j}(\bar{x}_{j})+\beta_{j}(P_{j},\bar{x}_{j})\geq 0,\;j\in J,j=N,$ (14i)
@@ IV-A Zero-Order Algorithm for Covariance Propagation
The zero-order algorithm of stochastic/robust NMPC [4, 5, 3] simplifies the Newton-type iteration for the OCP (14) by disregarding the sensitivities of the covariance propagation equations (14e) and (14f) and the sensitivities of the backoff terms $\beta(\cdot)$. The computational burden per Newton-type iteration is competitive with the nominal NMPC.
A limitation of the zero-order algorithm is that its convergence rate may slow down when the process noise covariance $W_{i}$ is substantial, as the backoff terms have a large effect on the convergence. We empirically found that for efficient optimization, $W_{i}$ should be carefully tuned to moderate the covariance propagation rather than accurately reflecting the process noise covariance of the actual system.
@@ IV-B Interior Point Method and Feedback Gain
We employ the nonlinear primal-dual interior point method to handle hard inequality constraints [15]. It should be noted that state-only inequality constraints need to be treated as soft constraints such as [16] to prevent the backoff term (III-C) from becoming excessively large, which leads to no feasible solution satisfying the tightened constraints (14h)–(14i) [4]. In practice, we also might have to clip the maximum values of the backoff terms to avoid ill-conditioned optimization problems.
@@ IV-C Algorithm Summary
6:     Propagate the state covariances by (14e) and (14f).
@@ V-A Full-Centroidal Formulations
Note that, although the reset map is the identity, the saltation matrix (II) is not the identity, and the guard saltation matrix (5) is not zero, due to the different values of $f_{m}^{-}$ and $f_{m+1}^{+}$. The guard condition consists of the distances between the swing feet and their targeted contact surfaces. Given a configuration $q$, the guard condition and its derivatives can be computed using forward kinematics and its Jacobian.
@@ V-B Comparison in Covariance Prediction
Fig. 3 illustrates the predicted base pose covariances using the three different methods. The state covariance is a $(12+n_{J})\times(12+n_{J})$ matrix, and the plots show the diagonal elements corresponding to the base pose dimensions. The comparison between methods (a) and (b) reveals that relying solely on the a priori covariance update (III-B1) leads to an overgrowth of the state covariance. This can lead to excessively large back-off terms (III-C) and (16), which in turn leads to a conservative solution or makes the numerical optimization challenging to solve. Conversely, the results from method (a) demonstrate that the combined use of a priori and a posteriori covariance updates effectively avoids the overgrowth of state covariance.
@@ V-C Simulation Comparison
Fig. 4 shows the success rates of the four methods in the two benchmarks. The proposed GS-SMPC outperformed the other methods in both scenarios, demonstrating its robustness against contact uncertainties. Figs. 57 display typical behaviors of Tachyon 3 and EVAL-03 during the experiments. As illustrated in Figs. 5 and 6, GS-SMPC successfully generated trajectories that avoided collisions while preventing joint torque saturation despite the discrepancies between expected and actual terrain heights. While SMPC also managed to avoid collisions, it encountered ill-conditioned optimization problems (e.g., too small step size in Fig. 6) caused by overly large backoff terms, as also illustrated in Fig. 3. Is should be noted that even though we attempted SMPC with wide range of $\sigma$ in (19), but could not find successfull one. HMPC and MPC were unable to accurately predict constraint margins, leading to intense collisions following the terrain height mispredictions. Fig. 7 also demonstrated effectiveness of GS-SMPC in preventing self-collisions while enhancing the unilateral force constraints under terrain height disturbances. The robot frequently experienced self-collisions between the left and right legs due to the contact disturbances when using the other MPC methods. However, as suggested by the success rate for EVAL-03 in Fig. 4, the proposed method still encountered challenges in preventing the robot from falling down, e.g., when a swing foot collided with bumps.
@@ V-D Hardware Experiments
Fig. 8 displays snapshots and plots of Tachyon 3’s perceptive locomotion using the proposed GS-SMPC. Tachyon 3 adeptly navigated over a step and managed to ascend and descend a series of two steps via GS-SMPC, which was implemented on the robot’s on-board PC. The average computational time of GS-SMPC was 29 ms while that of the nominal NMPC (i.e., NMPC without covariance propagation) was 26 ms measured on the on-board PC (CPU: Intel(R) Core(TM) i7-8850 H CPU @ 2.606 GHz). Thanks to the efficient algorithm of the stochastic NMPC outlined in Section IV, our GS-SMPC showed only a slight increase (approximately 3 ms in this instance) in computational time compared to the nominal NMPC.
@@ VI CONCLUSIONS
This paper introduced a stochastic/robust NMPC for robust model-based legged locomotion against contact uncertainties. We leveraged the saltation matrices and output-feedback MPC technique in forecasting the future state covariance, of which the performance had been shown in the numerical studies. Furthermore, the proposed method could be implemented on a on-board PC by utilizing the efficient zero-order stochastic/robust NMPC algorithm that fully leverages the Riccati recursion to obtain the feedback gains.